OpenStudy (anonymous):

Test has 25 questions. must have 75% correct. How do you figure out the final grade

OpenStudy (anonymous):

25*4=100 x*4=75 x=75/4 x=18.75 must have at least 19 questions correct

OpenStudy (anonymous):

To figure out the final grade u divide the number of correct questions by 25 and multiply by 100. 19/25 *100 = 76% whereas 18/25 is 72%.